# ESP-MedSAM: Efficient Self-Prompting SAM for Universal Domain-Generalized Medical Image Segmentation

### v4

@@ 4.3 Comparison on Domain-Generalized Medical Image Segmentation
Furthermore, we evaluate the generalization capabilities of our ESP-MedSAM architecture on unseen target domains, which is provided in Table 5. Within task-specific architectures, the large-capacity TransUNet [54] and multi-task learning MADGNet [7] show better performance than nnUNet [5]. On the contrary, our method achieves overwhelming performance on all unseen domains with a significant rise of 1.74%, 8.13%, 12.80%, 21.64%, 16.71% and 4.92% over MADGNet on the Dice metric. Compared to the lightweight SAMs, the proposed ESP-MedSAM tackle the challenge of segmentation mask generations in retinal vessels, ultrasound cancer and microscopic nuclei segmentation tasks, improving the Dice score by more than 50%. Remarkably, our ESP-MedSAM framework illustrates a lower HD distance, which provides a more precise localization for segmentation targets. Overall, these results demonstrate the superior generalization capabilities of the proposed ESP-MedSAM on unseen domains of different medical modalities while eliminating the demand for manual prompts.
4.4 Ablation Study
To investigate the effectiveness of MMDKD, SPPG and QDMD modules, we further conduct a comprehensive ablation study on unseen target domains of six medical imaging modalities, displayed in Table 6. By removing the devised modules from ESP-MedSAM, in $1^{st}$ row, the standard fine-tuned SAM [13] serves as the ablation baseline. By separately adding the MMDKD ($2^{nd}$ row), SPPG ($3^{rd}$ row) and QDMD ($4^{th}$ row), the performance is increased with the average Dice of all modalities gain of 9.32%, 5.47%, 2.79%, respectively. Particularly, the MMDKD strategy decreases 67.90% of the parameters compared to SAM. The introducing SPPG module ($3^{rd}$ row) significantly reduces HD by 66.19%, which is more efficient than manual point prompts. The $5^{th}$ row to the $7^{th}$ row indicate the compatibility between each module. Overall, these ablation experiments prove that the MMDKD strategy significantly decreases the computation costs of ESP-MedSAM. SPPG eliminates the requirement for manual annotations. QDMD further enhances the generalization capabilities of the model across diverse modalities.
4.5 Discussion
4.5.1 Effectiveness of Patch Prompt Learning
In this section, we delve into the rationale behind the adoption of the patch prompt learning strategy. Initially, we examine the efficiency of SPPG with different patch sizes on source domains. As illustrated in Fig. 4 (left), based on the average performance on six modalities, setting a relatively large patch size proves beneficial for the patch category task, albeit it provides less semantic information than mask prompts. Conversely, reducing the patch size increases the prediction complexity, consequently leading to more errors or noise. Additionally, we conduct a comparison of our devised patch prompt with point, box and Low-Resolution (LR) mask prompt modes in Fig. 4 (right). Considering that automatic prompt generation aims to learn the representation of human annotations, all prompt types utilize manual annotations (ground truth) as inputs for the experiment. Observations reveal that the manual patch prompt outperforms both the box and point prompts. Therefore, the learning patch prompt approach proves more efficient for guiding the segmentation decoding.
4.5.2 Significance of Decoupling Strategy
In the design of the MMDKD strategy, we adopt a decoupling strategy to provide an independent encoding workflow for each modality. To qualitatively evaluate the effectiveness of ESP-MedSAM in learning discriminative representations, we make a feature comparison via T-SNE with the baseline (using a full-parameter sharing strategy). As shown in Fig. 5, the features produced by our method exhibit significant discriminability over the six medical image modalities on unseen domains, which benefits the following decoder in terms of segmentation masks significantly, enhancing the universal and generalized capabilities of ESP-MedSAM.
@@ 4.5.2 Significance of Decoupling Strategy
Figure 4: Hyper-parameter analysis of patch size (left) and comparison of model performance based on different prompt types (right). Each prompt type utilises manual annotations (ground truth) as the input.
@@
Figure 5: Feature comparison via T-SNE between the baseline model and our ESP-MedSAM framework on unseen target domains.
5 Conclusion
@@
In this paper, we have proposed the ESP-MedSAM framework for universal medical image segmentation. Specifically, The MMDKD strategy has been introduced to distil the knowledge from the foundation model to a modality controller and a modality controller to create a semi-parameter sharing image encoder for discriminative feature generation. Then, SPPG has been devised to automatically produce a set of high-quality patch prompts for assisting segmentation decoding. Finally, QDMD has customized a specific segmentation workflow for each modality. Extensive experiments have demonstrated that ESP-MedSAM has lower computational complexity than the standard SAM and outperforms SOTA task-specific architectures and lightweight SAMs in diverse medical imaging segmentation tasks, exhibiting superior zero-shot generalization and universal capabilities.

### v5

@@ IV-C Comparison with State-of-the-arts on Unseen Domains
Furthermore, we evaluate the generalization capabilities of our De-LightSAM architecture on unseen target domains, which is provided in Table IV. Within task-specific architectures, the large-capacity TransUNet [55] and multi-task learning MADGNet [12] show better performance than nnUNet [7]. On the contrary, our method achieves overwhelming performance on all unseen domains with a significant rise of 1.74%, 8.13%, 12.80%, 21.64%, 16.71% and 4.92% over MADGNet on the Dice metric. Compared to the lightweight SAMs, the proposed De-LightSAM tackles the challenge of segmentation mask generations in retinal vessels, ultrasound cancer and microscopic nuclei segmentation tasks, improving the Dice score by more than 50%. Remarkably, our De-LightSAM framework illustrates a lower HD distance, which provides a more precise localization for segmentation targets. Furthermore, we present the computation costs of each framework in Table V. In high-resolution medical image segmentation tasks, we observe that traditional task-specific architectures demonstrate acceptable inference speed but higher model complexity. The recent lightweight SAMs suffer from slow inference speed and high latency in the automatic mask generation mode due to the inefficient sliding window algorithm. In contrast, our De-LightSAM displays remarkable complexity-speed trade-offs. Overall, these results demonstrate the superior generalization capabilities of our proposed De-LightSAM framework on unseen domains across different medical imaging modalities with high computation efficiency.
TABLE VIII: Comparison with auto-prompting methods on domain-generalized medical image segmentation: $\mathcal{S}\rightarrow\mathcal{T}$.
Methods Prompt Types Dice (Avg.) $\uparrow$ HD (Avg.) $\downarrow$
CellSAM [56] Box 74.51 149.73
UN-SAM [30] LRMask 73.87 145.59
UV-SAM [20] Box + LRMask 74.24 148.14
PromptNucSeg [21] Point 73.95 153.72
SP-SAM [57] Box + Point 73.39 157.63
De-LightSAM Patch 75.86 140.60
@@
Figure 6: Hyper-parameter analysis of patch size (left) and comparison of model performance based on different prompt types (right). Each prompt type utilizes manual annotations (ground truth) as the input.
@@
Figure 7: Feature comparison via T-SNE between the baseline model (left) and our De-LightSAM framework (right) on unseen target domains.
IV-D Ablation Study
To investigate the effectiveness of our proposed DC-Encoder $M_{1}$, SP-Generator $M_{2}$ and QM-Decoder $M_{3}$ modules, we further conduct a comprehensive ablation study on unseen target domains of six medical imaging modalities, displayed in Table VI. By removing the devised modules from De-LightSAM, in $1^{st}$ row, the standard fine-tuned SAM [16] serves as the ablation baseline. By separately adding the MDKD ($2^{nd}$ row), SP-Generator ($3^{rd}$ row) and QM-Decoder ($4^{th}$ row), the performance is increased with the average Dice of all modalities gain of 7.45%, 5.47%, 2.79%, respectively. Particularly, the MDKD strategy decreases 67.90% of the parameters compared to SAM. The introducing SP-Generator module ($3^{rd}$ row) significantly reduces HD by 66.19%, which is more efficient than manual point prompts. The $5^{th}$ row to the $7^{th}$ row indicates the compatibility between each module. On this basis, our complete De-LightSAM framework ($8^{th}$ row) simultaneously adopts all three components to achieve the best performance across all modalities.
Furthermore, we conduct a detailed analysis of our MDKD strategy, as shown in Table VII. We consider DC-Encoder trained from scratch as the baseline ($1^{st}$ Row), which achieves 66.17% Dice and 192.49 HD. The addition of conventional knowledge distillation ($2^{nd}$ Row) improves performance to 72.04% Dice with reduced HD of 165.23. However, our proposed MDKD strategy ($3^{rd}$ Row) demonstrates superior effectiveness, providing an additional 3.82% Dice improvement and 24.63 HD reduction, validating the effectiveness of our decoupled distillation approach in handling diverse medical imaging scenarios. Overall, these ablation experiments prove that the DC-Encoder strategy significantly decreases the computation costs of De-LightSAM while maintaining competitive accuracy. SP-Generator eliminates the requirement for manual annotations. QM-Decoder further enhances the generalization capabilities of the model across diverse modalities by preventing cross-modal interference. The MDKD strategy enables effective knowledge transfer from foundation models to lightweight architectures, improving the generalization capabilities of De-LightSAM.
IV-E Discussion
IV-E1 Effectiveness of Patch Prompt Learning
In this section, we delve into the rationale behind the adoption of the patch prompt learning strategy. Initially, we replace our SP-Generator module with other auto-prompting methods in De-LightSAM and make a comparison in Table VIII. It can be observed that Cell-SAM [56], UN-SAM[20] and PromptNucSeg [21] perform better than UN-SAM [30] and SP-SAM [57] but require more parameter consumptions as they utilize additional segmentors (e.g., Segformer [58]) or detectors (e.g., DETR [59]) to produce prompts. In contrast, our SP-Generator module eliminates these demands and outperforms state-of-the-arts by directly extracting high-quality patch prompts from the image embedding. In addition, we examine the efficiency of SP-Generator with different patch sizes on source domains. As illustrated in Fig. 6 (left), based on the average performance on six modalities, setting a relatively large patch size proves beneficial for the patch category task, albeit it provides less semantic information than mask prompts. Conversely, reducing the patch size increases the prediction complexity, consequently leading to more errors or noise. Additionally, we conduct a comparison of our devised patch prompt with point, box and Low-Resolution (LR) mask prompt modes in Fig. 6 (right). Considering that automatic prompt generation aims to learn the representation of human annotations, all prompt types utilize manual annotations (ground truth) as inputs for the experiment. Observations reveal that the manual patch prompt outperforms both the box and point prompts. Therefore, the learning patch prompt approach proves more efficient for guiding the segmentation decoding.
IV-E2 Significance of Modality-Decoupled Framework
In the design of our MDKD strategy and image encoder, we adopt a decoupling strategy to provide an independent encoding workflow for each modality. To qualitatively evaluate the effectiveness of De-LightSAM in learning discriminative representations, we make a feature comparison via T-SNE with the baseline (using a full-knowledge sharing strategy). As shown in Fig. 7, the features produced by our method exhibit significant discriminability over the six medical image modalities on unseen domains, which benefits the following decoder in terms of segmentation masks significantly, enhancing the universal and generalized capabilities of De-LightSAM.
V Conclusion
@@
In this paper, we have proposed the De-LightSAM framework for universal medical image segmentation. Specifically, the DC-Encoder has been introduced with a modality controller to generate discriminative features for diverse medical modalities while maintaining computational efficiency. Then, SP-Generator has been designed to automatically produce high-quality dense prompt embeddings for guiding segmentation decoding, eliminating the dependency on manual annotations. Moreover, QM-Decoder has established independent decoding channels for each modality, preventing cross-modal knowledge interference and enhancing modality-specific performance. Finally, the MDKD strategy has been devised to distill both common and domain-specific knowledge from foundation models to our lightweight architecture, unleashing generalization potentials. Extensive experiments have demonstrated that De-LightSAM achieves significantly lower computational complexity than the standard SAM and outperforms state-of-the-arts across diverse medical imaging segmentation tasks, exhibiting superior domain generalization capabilities.